An equipment purchased at a cost $80,000 by a local company is being depreciated using MACRS method as a 5-year property. At the end of four years, the management decided to sell the equipment for a modest price of $20,000. The company is in the 34% tax bracket. Compute the tax consequence on the sale of this equipment.

A.

$6,800

B.

$926.40

C.

$533.12

D.

None of these
Answer #1

The depreciation for the first four years under MACRS

= 20%+32%+19.20%+11.52% = 82.72%

Book value = 80000-(82.72% of 80000) = 80000-66176 = 13824

The gain on the sale = 20000-13824 = 6176

The tax consequence is 34% of the 6176 = 2100

option(D) none of these